OpenStudy (anonymous):

Okay guise i need your utmost help i have a test tomorrow on this and on how to graph it, I am having trouble finding the derivative of -2x/((x^2-1)^2) and then on how to graph it :( help!

OpenStudy (dumbcow):

looks like quotient rule: = (f'g - fg')/g^2 f = -2x f' = -2 g = (x^2-1)^2 For g' use chain rule u = x^2 -1 du = 2x g' = d/du (u^2) *du g' = 4x(x^2 -1) = 4x^3-4x (f/g)' = (-2(x^2-1)^2 - (-2x)(4x^3-4x))/ (x^2-1)^4 then expand and simplify = (6x^4-4x^2-2)/(x^2-1)^4

OpenStudy (dumbcow):

as far as graphing it. do you need to graph the original function or the derivative? either way find the vertical and horizontal asymptotes, then plug in points

OpenStudy (anonymous):

well i found the first derivative from the original equation and got my asymptotes, So my question now is do I plug in points to the original equation?

OpenStudy (dumbcow):

it depends on which function you are graphing

OpenStudy (anonymous):

it tells me to sketch the graph of f , meaning the original given equation so that means i plug in points there?

OpenStudy (dumbcow):

ok then yes, pick x-values on either side of vertical asymptotes and plug them into original equation

OpenStudy (anonymous):

okay thank you! this will help me alot on my test tomorrow :)